# Geometric quantum thermodynamics: A fibre bundle approach

### v2

@@ IV Geometric theory of quantum thermodynamics
We begin by addressing what might, at first, seem to be an inconsistency. The theory developed in Refs. [20, 21] presents a gauge theory of the thermodynamic group $\mathrm{G}_{T}$. On the other hand, the potential $i\dot{u}_{t}u^{\dagger}_{t}$, defined for the covariant derivative, takes values in the Lie algebra of the unitary group $\mathrm{U}(d)$, and not in that of the thermodynamic group. We will shortly see that this is not an inconsistency and that we are actually dealing with two distinct (although related) geometrical structures, and not just one.
@@
Moreover, it is important to note that, since $\mathrm{G}_{T}$ is not a fundamental gauge group, such as the ones in the standard model, but instead emerges from a lack of information about the system at hand, we should not expect it to behave in the same way as the group $\mathrm{U}(1)$ does in electrodynamics, for example. Indeed, as will become clear later on, this emergent thermodynamic gauge has a rich geometric structure that sets it apart from those commonly studied in field theories.
@@
To construct the relevant fibre bundles for the theory, we note that since all quantities of interest, like the density operator $\rho_{t}$ and the Hamiltonian $H_{t}$, depend only on time. Therefore, by analogy with the gauge theories of the standard model, the base space of our bundle should be time, which can be taken to be the real line $\mathds{R}$. Since $\mathds{R}$ is contractible [10], our bundle is necessarily trivial by Theorem 1 in Appendix A. As such, we take the liberty to employ the isomorphism between a trivial bundle and a product bundle of the base space and the fibre, and we always regard quantum thermodynamics as taking place on the product bundle.
@@
For the fibre, seeing the definition of the covariant derivative, we take the entire unitary group $\mathrm{U}(d)$ acting on the Hilbert space of dimension $d$. We further note that, since $\mathrm{G}_{T}$ is a subgroup of $\mathrm{U}(d)$, any structure regarding the thermodynamic group can be regarded as a substructure of this principal $\mathrm{U}(d)$-bundle.
@@
Thus, the first geometrical setup for quantum thermodynamics is a principal $\mathrm{U}(d)$ bundle $\xi=(\mathds{R}\times\mathrm{U}(d),\pi,\mathrm{U}(d),\mathds{R})$, with $\pi$ having the same meaning as before. We note that it makes sense for the gauge group to be $\mathrm{U}(d)$, as we are not currently concerned with any specific measurements of thermodynamic quantities that would require gauge invariance under the action of $\mathrm{G}_{T}$. Since this bundle arises from the unitary matrices $u_{t}$ that diagonalize the Hamiltonian $H_{t}$, we can observe that the gauge freedom associated with the unitary group is that of a choice of basis, which we use to work in the eigenbasis of the Hamiltonian.
@@
The next step in characterising the bundle’s geometry is verifying that the potential present in the covariant derivative defines a connection form on $\xi$. This can be seen to be true by noting that if we slightly change the definition of the potential from $i\dot{u}_{t}u_{t}^{\dagger}$ to $A_{t}=\dot{u}_{t}u_{t}^{\dagger}$ (thus changing the covariant derivative to $\mathrm{D}_{t}\cdot=\partial_{t}+[A_{t},\cdot]$), we see that $A_{t}$ is precisely a right-invariant Maurer-Cartan form on $\mathrm{U}(d)$. Thus, by Theorem 2 in Appendix A, $A_{t}$ defines a connection on $\xi$.
@@
We take $\mathds{V}$ to be the vector space of $d\times d$ Hermitian matrices and have the group $\mathrm{U}(d)$ act on it via the adjoint representation, such that $u\cdot\nu=u\nu u^{\dagger}$, where $u\in\mathrm{U}(d)$ and $\nu\in\mathds{V}$. Since our original principal bundle is trivial, we can always take the associated bundle to be the product vector bundle with total space $\mathds{R}\times\mathds{V}$. This means that each point on $\eta$ is a tuple $(t,\nu)$, where $t$ is an instance of time and $\nu$ is some $d\times d$ Hermitian matrix. By analogy with the notion of matter fields as sections of the associated bundle discussed earlier, and noting that a global section of $\eta$ associates each point $t\in\mathds{R}$ with a Hermitian matrix $\nu\in\mathds{V}$, we see that the "matter fields" of the gauge theory of quantum thermodynamics are simply time-dependent Hermitian observables.
Given the above argument, we naturally consider the density matrices not as emergent gauge potentials, as proposed in Refs. [20, 21], but as specific matter fields of the theory. Thus, in this framework, the thermodynamic quantities of interest arise as functionals of the matter fields (observables), much like in a quantum field theory where several physical properties arise as functionals of matter fields (for example, Dirac or Klein-Gordon fields) [8, 9, 25].
@@
We now return to the topic of the gauge symmetry associated with the thermodynamic group $\mathrm{G}_{T}$. As defined in Eq. (3), it is more accurately described as a time-dependent family of Lie groups. Thus, for different points in the base space $\mathds{R}$, $\mathrm{G}_{T}$ is not necessarily the same group. As such, we do not, in general, have a single principal $\mathrm{G}_{T}$-bundle that can be defined as a subbundle of $\xi$ with base space $\mathds{R}$. We can, however, note that at any instant $t$, we can form a trivial principal bundle with total space $\{t\}\times\mathrm{G}_{T}^{t}$, where by $\mathrm{G}_{T}^{t}$ we denote the thermodynamic group at time $t$. This bundle has a connection defined by the Maurer-Cartan form of $\mathrm{G}_{T}^{t}$, similarly to how we define the connection for the $\mathrm{U}(d)$ bundle. It is this structure that carries along the emergent gauge symmetry related to the redundant information contained in the density matrix. This can be seen by noting that our construction of the associated bundle applies identically in this case.
This second geometric structure defined above formalises the idea that the emergent gauge of quantum thermodynamics arises only concerning measurements, which we assume are taken at some specific time $t$. This can be understood as the geometric meaning of the coarse-graining introduced in Ref. [20]. In a situation where the degeneracies of the Hamiltonian (and consequently the thermodynamic group) do not change continuously in time but vary in specific times $\{t_{1},t_{2},...,t_{m}\}$, we can define, instead of bundles over points, bundles over time intervals $0,t_{1}),[t_{1},t_{2}),...[t_{m-1},t_{m}),[t_{m},\infty)$ equipped with connections given by the Maurer-Cartan forms and Theorem 2. This construction may be useful in a geometrical study of quench dynamics, where additional interaction terms in the Hamiltonian are turned on at a discrete set of times.
@@
In an effort to make the discussion above easier to digest, we again evoke the example of electrodynamics for a final comparison, as shown in Table 1.
@@
V Discussion
@@
In this work, we have mathematically formalized the gauge theory of quantum thermodynamics, whose main ideas were put forward in Refs. [20, 21]. Specifically, the theory was formulated through the geometry of a principal $\mathrm{U}(d)$-bundle with an additional geometric structure associated with measurements (namely, the thermodynamic group $\mathrm{G}_{T}$).
@@
In our argument, we have separated which mathematical objects belong to each geometric structure associated with the theory. Furthermore, by constructing associated vector bundles, we were able to create an analogy between gauge-theoretic matter fields and time-dependent Hermitian observables in quantum thermodynamics. This result helps strengthen the link between the theory proposed in Refs. [20, 21] and the gauge theories in fundamental physics, such as Yang-Mills theory. In other words, we put thermodynamics in the same mathematical language as the fundamental theories in physics.
@@
Furthermore, although the structure associated with the thermodynamic group does not, in general, form a single principal fibre bundle, it may form a collection of them if the degeneracies of the observables under consideration change discretely over time. In particular, formalizing these ideas geometrically allows us to propose a series of questions regarding how the topological properties of the thermodynamic group influence the thermodynamics of a quantum system.
@@
Another perhaps less obvious direction is the study of relative homology and cohomology [10]. Since the thermodynamic group is a Lie subgroup of the unitary group $\mathrm{U}(d)$, one may construct relative homologies and cohomologies between them. This may be of interest, as both $\mathrm{G}_{T}$ and $\mathrm{U}(d)$ are relevant gauge groups of the theory. In particular, since relative Rham cohomology relates to differential forms on $\mathrm{U}(d)$ that vanish on $\mathrm{G}_{T}$, and gauge-invariant quantities in this framework carry integrals over the thermodynamic group, it makes sense that it may have some meaning in the theory.
Additionally, we know that in the theory of classical thermodynamics, there are many important inequalities, such as inequalities constraining changes in the free energies of a system [1]. Meanwhile, in topology, there are known relations between the critical points of functions and the topology of their domain. Namely, Morse theory [10] does so by constructing inequalities that relate the numbers of critical points with different indices of the function to the Betti numbers (ranks of homology and cohomology groups) of the space. From this perspective, we believe that Morse theory may offer insight into quantum thermodynamic equilibrium by examining the critical points of certain properly defined free energies. We also believe that it may be possible for a generalised Clausius inequality, such as the one hypothesised in [21], to have some relation to the structure of the thermodynamic group, whether through its algebraic or geometric/topological properties.
One last, perhaps less speculative direction to follow relates to the study of the thermodynamic group in the context of Lie group bouquets [27]. As discussed earlier, due to its time dependence, the thermodynamic group by itself doesn’t form a proper fibre bundle for all times (and thus doesn’t define a gauge theory). The concept of a Lie Group Bouquet, introduced in the seminal 1966 paper by Douady and Lazard [27], formalises the notion of a family of Lie groups parametrised by some other manifold (which, in this case, would be $\mathds{R}$). It is immediately obvious that the thermodynamic group forms one such bouquet, and it should be worthwhile to study it as such.
VI Acknowledgments

### v3

@@ IV Geometric theory of quantum thermodynamics
We begin by addressing what might, at first, seem to be an inconsistency. The theory developed in [20, 21] presents a gauge theory of the thermodynamic group $\mathrm{G}_{T}$. On the other hand, the potential $i\dot{u}_{t}u^{\dagger}_{t}$, defined for the covariant derivative, takes values in the Lie algebra of the unitary group $\mathrm{U}(d)$ and not in that of the thermodynamic group. Here, $\mathrm{U}(d)$ denotes the unitary group acting on the full system Hilbert space of dimension $d$, as defined in Sec. II. We will shortly see that this is not an inconsistency and that we are actually dealing with two distinct (although related) geometrical structures, and not just one.
@@
Moreover, it is important to note that since $\mathrm{G}_{T}$ is not a fundamental gauge group, such as the ones in the standard model, but instead emerges from a lack of information about the system at hand, we should not expect it to behave in the same way as the group $\mathrm{U}(1)$ does in electrodynamics, for example. Indeed, as will become clear later on, this emergent thermodynamic gauge has a rich geometric structure that sets it apart from those commonly studied in field theories.
@@
To construct the relevant fiber bundles for the theory, we note that all quantities of interest, such as the density operator $\rho_{t}$ and the Hamiltonian $H_{t}$, depend only on time. Therefore, by analogy with the gauge theories of the standard model, the base space of our bundle should be time, which can be taken to be the real line $\mathds{R}$. Since $\mathds{R}$ is contractible [10], our bundle is necessarily trivial by Theorem 1 in Appendix A. As such, we take the liberty to employ the isomorphism between a trivial bundle and a product bundle of the base space and the fiber, and we always regard quantum thermodynamics as taking place on the product bundle.
@@
For the fiber, seeing the definition of the covariant derivative, we take the entire unitary group $\mathrm{U}(d)$ acting on the Hilbert space of dimension $d$. We further note that since $\mathrm{G}_{T}$ is a subgroup of $\mathrm{U}(d)$, any structure regarding the thermodynamic group can be considered a substructure of this principal $\mathrm{U}(d)$-bundle.
@@
Thus, the first geometrical setup for quantum thermodynamics is a principal $\mathrm{U}(d)$ bundle $\xi=(\mathds{R}\times\mathrm{U}(d),\pi,\mathrm{U}(d),\mathds{R})$, with $\pi$ having the same meaning as before. We note that it makes sense for the gauge group to be $\mathrm{U}(d)$, as we are currently not concerned with any specific measurements of thermodynamic quantities that would require gauge invariance under the action of $\mathrm{G}_{T}$. Since this bundle arises from the unitary matrices $u_{t}$ that diagonalize the Hamiltonian $H_{t}$, we can observe that the gauge freedom associated with the unitary group is that of a choice of basis, which we use to work in the eigenbasis of the Hamiltonian.
@@
The next step in characterizing the geometry of the bundle is verifying that the potential present in the covariant derivative defines a form of connection on $\xi$. This can be seen to be true by noting that if we slightly change the definition of the potential from $i\dot{u}_{t}u_{t}^{\dagger}$ to $A_{t}=\dot{u}_{t}u_{t}^{\dagger}$ (thus changing the covariant derivative to $\mathrm{D}_{t}\cdot=\partial_{t}+[A_{t},\cdot]$), we see that $A_{t}$ is precisely a right-invariant Maurer-Cartan form on $\mathrm{U}(d)$. Thus, by Theorem 2 in Appendix A, $A_{t}$ defines a connection on $\xi$.
@@
We take $\mathds{V}$ to be the vector space of $d\times d$ Hermitian matrices and have the group $\mathrm{U}(d)$ act on it through the adjoint representation such that $u\cdot\nu=u\nu u^{\dagger}$, where $u\in\mathrm{U}(d)$ and $\nu\in\mathds{V}$. Since our original principal bundle is trivial, we can always take the associated bundle to be the product vector bundle with total space $\mathds{R}\times\mathds{V}$. This means that each point on $\eta$ is a tuple $(t,\nu)$, where $t$ is an instance of time and $\nu$ is some $d\times d$ Hermitian matrix. By analogy with the notion of matter fields as sections of the associated bundle discussed earlier, and noting that a global section of $\eta$ associates each point $t\in\mathds{R}$ with a Hermitian matrix $\nu\in\mathds{V}$, we see that the "matter fields" of the gauge theory of quantum thermodynamics are simply time-dependent Hermitian observables.
Given the above argument, we naturally consider the density matrices not as emergent gauge potentials, as proposed in [20, 21], but as specific matter fields of the theory. Thus, in this framework, the thermodynamic quantities of interest arise as functionals of the matter fields (observables), much like in a quantum field theory where several physical properties arise as functionals of matter fields (for example, Dirac or Klein-Gordon fields) [8, 9, 25].
@@
We now return to the topic of the gauge symmetry associated with the thermodynamic group $\mathrm{G}_{T}$. As defined in Eq. (3), it is more accurately described as a time-dependent family of Lie groups. Thus, for different points in the base space $\mathds{R}$, $\mathrm{G}_{T}$ is not necessarily the same group. As such, we do not, in general, have a single principal $\mathrm{G}_{T}$-bundle that can be defined as a sub-bundle of $\xi$ with base space $\mathds{R}$. However, we can note that at any instant $t$, we can form a trivial principal bundle with total space $\{t\}\times\mathrm{G}_{T}^{t}$, where by $\mathrm{G}_{T}^{t}$ we denote the thermodynamic group at time $t$. This bundle has a connection defined by the Maurer-Cartan form of $\mathrm{G}_{T}^{t}$, similarly to how we define the connection for the bundle $\mathrm{U}(d)$. It is this structure that carries along the emergent gauge symmetry related to the redundant information contained in the density matrix. This can be seen by noting that our construction of the associated bundle applies identically in this case.
This second geometric structure defined above formalizes the idea that the emergent gauge of quantum thermodynamics arises only concerning measurements, which we assume are taken at some specific time $t$. This can be understood as the geometric meaning of the coarse-graining introduced in Ref. [20]. In a situation where the degeneracies of the Hamiltonian (and consequently the thermodynamic group) do not change continuously in time but vary in specific times $\{t_{1},t_{2},...,t_{m}\}$, we can define, instead of bundles over points, bundles over time intervals $0,t_{1}),[t_{1},t_{2}),...[t_{m-1},t_{m}),[t_{m},\infty)$ equipped with connections given by the Maurer-Cartan forms and Theorem 2. This construction may be useful in a geometrical study of quench dynamics, where additional interaction terms in the Hamiltonian are turned on at a discrete set of times.
@@
The vanishing of the curvature should not be interpreted as a lack of physical content, but rather as a direct consequence of the operational structure of thermodynamics. In the present framework, the base manifold is identified with physical time, reflecting the fact that thermodynamic processes are defined through temporal protocols —preparations, drivings, and measurements— rather than through spatially extended gauge fields. As a result, the associated gauge structure organizes how thermodynamic descriptions are compared along time, rather than encoding dynamical forces in the sense of Yang–Mills theories.
Despite the absence of curvature, the thermodynamic connection plays a nontrivial physical role. It defines a covariant notion of temporal transport for thermodynamic quantities, ensuring consistency under time-dependent changes of description induced by restricted measurements. In particular, when the density matrix is interpreted as a section of an associated bundle, the connection specifies how states at different times are compared in a gauge-invariant manner. This covariant transport is essential for the definition of physically meaningful notions of work, heat, and entropy.
Importantly, the physical content of the theory resides in the global properties of the connection along finite-time processes rather than in local curvature. Even in one dimension, the connection, paired with the coarse-graining, can generate nontrivial holonomies that encode the cumulative effect of gauge transformations along a thermodynamic protocol. In this sense, irreversibility and path dependence arise not from curvature, but from the structure of time-ordered transport in the space of thermodynamic descriptions.
In an effort to make the discussion above easier to digest, we again invoke the electrodynamic example for a final comparison, as shown in Table 1.
@@
The thermodynamic gauge group $\mathrm{G}_{T}$ is, in general, time dependent. This time dependence is not a formal artifact, but reflects a physically meaningful feature of thermodynamic processes: the information that is accessible to an observer may change during the evolution. In the present framework, the gauge group is determined by the degeneracy structure of the observables that are monitored during the process. Whenever this degeneracy structure changes in time, the associated thermodynamic gauge symmetry must also change.
Such situations naturally arise in driven quantum systems, where external control parameters modify the Hamiltonian and, consequently, the spectrum and its degeneracies. Examples include quenches, slow parameter ramps, or protocols crossing level crossings or symmetry-breaking points. In these cases, the equivalence classes of states defined by thermodynamic coarse-graining are themselves time dependent, and the appropriate description is not a single fixed gauge group acting globally in time, but a family of thermodynamic groups $\mathrm{G}_{T}$ acting locally along the process.
From a geometric perspective, this implies that the thermodynamic bundle is defined piecewise in time, with possibly distinct gauge structures over different temporal intervals. The connection introduced in this work provides a consistent way to relate thermodynamic descriptions across these intervals, ensuring gauge-covariant transport even when the underlying symmetry group changes. In this sense, the time dependence of $\mathrm{G}_{T}$ encodes changes in the informational constraints imposed by the measurement protocol, rather than changes in the microscopic dynamics itself.
This structure has direct physical implications. Since gauge invariance determines which quantities are thermodynamically meaningful, changes in $\mathrm{G}_{T}$ modify the set of admissible gauge-invariant quantities during the process. As a result, the classification of energy exchanges into work, heat, and entropy becomes explicitly protocol dependent, reflecting the evolving accessibility of information. The present framework thus provides a geometric language to describe nonequilibrium thermodynamic processes in which the notion of thermodynamic equivalence evolves in time.
V Application: The LMG model
To illustrate how the geometric thermodynamic framework developed in this work applies to a concrete physical system, we consider the Lipkin-Meshkov-Glick (LMG) model [27, 28, 29]. This model provides a particularly transparent example, as its symmetry properties and degeneracy structure are well understood and can be externally controlled through time-dependent parameters. Here we present a description of this model and relate it to the geometric structure constructed here. An extensive numerical analysis of this model was presented in the context of the gauge theory of thermodynamics in [21], where the invariant work, heat, and entropy were calculated for quench protocols.
The LMG model describes a system of $N$ spin-$\tfrac{1}{2}$ particles with infinite-range interactions and is conveniently expressed in terms of collective spin operators $J_{\alpha}=\tfrac{1}{2}\sum_{i=1}^{N}\sigma_{i}^{\alpha}$, where $\alpha=x,y,z$. A commonly used form of the Hamiltonian reads
$H_{t}=-\frac{\lambda_{t}}{N}\left(J_{x}^{2}+\gamma J_{y}^{2}\right)-h_{t}\,J_{z},$ (12)
where $\lambda_{t}$ is an interaction strength, $\gamma$ is an anisotropy parameter, and $h_{t}$ is a transverse field. Both $\lambda_{t}$ and $h_{t}$ can be controlled externally and vary over time according to a prescribed protocol.
For fixed parameters, the Hamiltonian is invariant under permutations of spins and conserves the total angular momentum. As a consequence, the Hilbert space decomposes into invariant subspaces labeled by the total spin quantum number $J$, and the energy spectrum generally exhibits degeneracies associated with this symmetry. These degeneracies define the equivalence classes of states under thermodynamic coarse graining and, in the present framework, determine the thermodynamic gauge group $\mathrm{G}_{T}$ acting on the associated bundle of density operators.
When the system is driven by a time-dependent protocol, the degeneracy structure of the Hamiltonian may change. For example, variations in the transverse field $h_{t}$ or the interaction strength $\lambda_{t}$ can lift or create degeneracies, particularly near critical points that separate distinct phases of the model. As a result, the thermodynamic gauge group becomes explicitly time dependent, $\mathrm{G}_{T}\rightarrow\mathrm{G}_{T}(t)$, reflecting the evolving symmetry and information content of the system.
Within the geometric formulation developed in this work, this situation is naturally described by a family of thermodynamic gauge groups defined along the time axis. The density matrix is interpreted as a section of an associated bundle, and the thermodynamic connection ensures gauge-covariant transport of states under time evolution, even when the underlying gauge group changes. Importantly, although the base manifold is one-dimensional and the curvature vanishes identically, the connection remains nontrivial and encodes the physically relevant parallel transport between thermodynamic descriptions defined at different times.
The physical implications of this structure are directly observable at the level of thermodynamic quantities. Since gauge invariance determines which observables are thermodynamically significant, changes in $\mathrm{G}_{T}$ modify the set of admissible gauge-invariant quantities during the protocol. In particular, the decomposition of energy variations into work and heat becomes explicitly dependent on the evolving thermodynamic symmetry, reflecting changes in the informational constraints imposed by the measurement scheme. This provides a geometric explanation for the protocol dependence of thermodynamic quantities in driven many-body systems.
The LMG model thus offers a concrete realization of the central ideas of this work. It shows how changes in degeneracy and symmetry naturally induce a time-dependent thermodynamic gauge structure and how the associated bundle formulation provides a consistent and physically transparent description of nonequilibrium thermodynamic processes. Although no explicit dynamics is required for the present discussion, this example demonstrates that the refined geometric framework developed here extends beyond a purely formal construction and captures physically relevant features of realistic many-body systems.
The time dependence of the thermodynamic gauge group $\mathrm{G}_{T}(t)$ has direct consequences for the definition and interpretation of thermodynamic quantities such as work, heat, and entropy. In the present framework, these quantities are defined through gauge-covariant variations of the density matrix along the time axis. As a result, their physical meaning is intrinsically tied to the connection on the thermodynamic bundle.
For a fixed thermodynamic gauge group, the infinitesimal change of the internal energy,
[…]
admits a gauge-invariant decomposition into work and heat, with the first term identified as work and the second as heat. This separation relies on the assumption that the notion of thermodynamic equivalence, encoded in $\mathrm{G}_{T}$, remains unchanged during the process.
When the thermodynamic gauge group becomes time-dependent, this assumption no longer holds globally. Changes in the degeneracy structure of the Hamiltonian modify the equivalence classes of states under thermodynamic coarse-graining and therefore alter which variations of the density matrix are physically distinguishable. In geometric terms, the subspaces defined by the connection are themselves time dependent.
[…]
This effect has a clear physical interpretation in the context of the LMG model. When a time-dependent protocol lifts or creates degeneracies in the energy spectrum, previously indistinguishable microstates become distinguishable, or vise versa. This change modifies the entropy associated with coarse-grained descriptions, even in the absence of energy exchange [21]. The resulting entropy variation is therefore geometric in origin, arising from parallel transport between thermodynamic descriptions defined by different gauge groups.
From this perspective, entropy production receives a contribution associated with the mismatch between thermodynamic descriptions at different times, encoded in the holonomy generated by the thermodynamic connection and coarse-graining. Although the curvature vanishes identically due to the one-dimensional base manifold, the connection still generates nontrivial parallel transport, which captures the cumulative effect of changing informational constraints along the process.
The LMG model thus illustrates how the refined geometric structure developed in this work provides a unified description of energetic and informational contributions to thermodynamic quantities. Work and heat are no longer determined solely by microscopic dynamics, but also by the evolution of the thermodynamic gauge structure. This demonstrates that the connection introduced here is not merely a mathematical artifact but a physically meaningful object encoding how thermodynamic quantities transform under time-dependent coarse-graining and symmetry changes.
VI Discussion
@@
In this work, we have mathematically formalized the gauge theory of quantum thermodynamics, whose main ideas were put forward in Refs. [20, 21]. Specifically, the theory was formulated through the geometry of a principal $\mathrm{U}(d)$-bundle with an additional geometric structure associated with measurements (namely the thermodynamic group $\mathrm{G}_{T}$).
@@
In our argument, we have separated which mathematical objects belong to each geometric structure associated with the theory. Furthermore, by constructing associated vector bundles, we were able to create an analogy between gauge-theoretic matter fields and time-dependent Hermitian observables in quantum thermodynamics. This result helps to strengthen the link between the theory proposed in Refs. [20, 21] and the gauge theories in fundamental physics, such as Yang-Mills theory. In other words, we put thermodynamics in the same mathematical language as fundamental theories in physics.
@@
Furthermore, although the structure associated with the thermodynamic group does not, in general, form a single principal fiber bundle, it may form a collection of them if the degeneracies of the observables under consideration change discretely over time. In particular, formalizing these ideas geometrically allows us to propose a series of questions regarding how the topological properties of the thermodynamic group influence the thermodynamics of a quantum system.
Finally, we note that thermodynamic quantities depend only on energy differences and are therefore invariant under global shifts of the Hamiltonian by a multiple of the identity. Within the present framework, such shifts correspond to transformations that leave all gauge-invariant thermodynamic observables unchanged. This reflects the fact that the thermodynamic gauge structure encodes redundancies associated with informational redundant degrees of freedom in the description, including both coherences within degenerate subspaces and global energy offsets. The proposed gauge formulation thus naturally incorporates the fundamental insensitivity of thermodynamics to the choice of energy zero.
@@
Another perhaps less obvious direction is the study of relative homology and cohomology [10]. Since the thermodynamic group is a Lie subgroup of the unitary group $\mathrm{U}(d)$, relative homologies and cohomologies between them can be constructed. This may be of interest, as both $\mathrm{G}_{T}$ and $\mathrm{U}(d)$ are relevant gauge groups of the theory. In particular, since relative Rham cohomology relates to differential forms in $\mathrm{U}(d)$ that vanish in $\mathrm{G}_{T}$, and gauge-invariant quantities in this framework carry integrals over the thermodynamic group, it makes sense that it may have some meaning in theory.
Additionally, we know that in the theory of classical thermodynamics, there are many important inequalities, such as inequalities that restrict changes in the free energies of a system [1]. Meanwhile, in topology, there are known relations between the critical points of functions and the topology of their domain. That is, Morse theory [10] does so by constructing inequalities that relate the numbers of critical points with different indices of the function to the Betti numbers (ranks of homology and cohomology groups) of the space. From this perspective, we believe that Morse theory may offer insight into quantum thermodynamic equilibrium by examining the critical points of certain properly defined free energies. In this context, it may be possible to derive a generalized Clausius inequality, such as the one hypothesized in [21], that is related to the structure of the thermodynamic group through its algebraic or geometric/topological properties [34].
The last, perhaps less speculative direction to follow relates to the study of the thermodynamic group in the context of Lie group bouquets [30]. As discussed earlier, due to its time dependence, the thermodynamic group by itself does not form a proper fiber bundle for all times (and thus does not define a gauge theory). The concept of a Lie Group Bouquet, introduced in the seminal 1966 paper by Douady and Lazard [30], formalizes the notion of a family of Lie groups parametrized by some other manifold (which, in this case, would be $\mathds{R}$). It is immediately obvious that the thermodynamic group forms one such bouquet, and it should be worthwhile to study it as such.
VII Acknowledgments